OpenStudy (anonymous):

an electron moving with velocity c/2 having de brogle eqal to wavelength of photon.the ratio of their energy is????????????

OpenStudy (anonymous):

If you post it here http://openstudy.com/groups/mathematics#/groups/mathematics People WILL help you. Even though its not math. Promise. I do it ALLLLLLLLL the time ^-^

OpenStudy (anonymous):

Since the electron is moving at relativistic velocities, we must use the relativistic equation for the total energy of a particle. This is\[E=\sqrt{p^2c^2+(mc^2)^2}\]where \(p=mv\) is the momentum of the particle of mass \(m\) and velocity \(v\). We are told that the velocity of the electron is \(c/2\), where \(c\) is teh speed of light. Substituting in we get \[E=\sqrt{\left(\frac{mc}{2}\right)^2c^2+(mc^2)^2}\]which simplifies down to \[E=\sqrt{\left(\frac{m^2c^4}{4}\right)+(m^2c^4)}=\sqrt{\left(\frac{5}{4}\right)m^2c^4}=\frac{\sqrt{5}}{2}mc^2\] Now a particle with deBroglie wavelength \(\lambda\) will have an energy of \[E=\frac{hc}{\lambda}\]where \(h\) is Planck's Constant. We now equate the two expressions to give \[\frac{hc}{\lambda}=\frac{\sqrt{5}}{2}mc^2\] Rearranging to make \(\lambda\) the subject we get \[\lambda=\frac{2h}{\sqrt{5}mc}\]and hence if the mass of the electron is \(m=9.109\times10^{-31}\) kg, then the deBroglie wavelength is thus \[\lambda=\frac{2(6.626\times10^{-34})}{\sqrt{5}(9.109\times10^{-31})(2.998\times10^8)}=2.170\times10^{-12}\rm{m}\] What this means is that one can use relativistic electrons for imaging purposes as if they had a photon wavelength of \(2.170\times10^{-12}\) m, and so can resolve smaller features than is capable using visible wavelengths (around 500 nm).

OpenStudy (anonymous):

i guess johnyymca is correct